OpenStudy (anonymous):

0.2(x+3)-4(2x+3)=3.4

OpenStudy (anonymous):

distribute and combine like terms...check that you wrote the 2nd parenthesis set correct

OpenStudy (anonymous):

0.2(x+3) - 4(2x-3) = 3.4 0.2x + 0.6 - 8x + 12 = 3.4 -7.8x + 12.6 = 3.4 12.6 = 3.4 + 7.8x 9.2 = 7.8x x = 9.2/7.8 = 92/78 = 46/39